Solve the formula for R. You don't have to "simplify" anything.
H=sqrt(2R+1)-3
The 2R+1 is under the sqrt symbol in the question.

Answer by Boreal(15235)

H+3=sqrt(2R+1)
square both sides
H^2+6H+9=2R+1
R=(H^2+6H+8)/2
=============
check
let r=4
then h=sqrt(9)-3 or 0
Therefore R=8/2=4